OpenStudy (anonymous):

Help with the steps for ax+2xy=14 solve for y

OpenStudy (anonymous):

ax + 2xy = 14 x(a + 2y) = 14 /x /x ---------------- a + 2y = 14/x -a -a ---------------- 2y = 14/x - a /2 /2 /2 --------------- y = 14/2x - a/2

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem, if I helped you do you mind giving me a medal? thanks!